the sum of two odd integers is 11 times more than three times the larger number. find the integers
Answer by macston(5194)

Using consecutive odd integers and editing "11 times more" to "11 more":
(Using original wording, there is no solution)
x=smaller integer; x+2=larger integer
x+(x+2)=3(x+2)-11
2x+2=3x+6-11 Add 5 to each side.
2x+7=3x Subtract 2x from each side.
7=x ANSWER : The smaller integer is 7
x+2=7+2=9 ANSWER 2: The larger integer is 9